OpenStudy (anonymous):

how do i divide -2/3 by -18?

OpenStudy (vishweshshrimali5):

ok so lets start from here \[\frac{-ve}{-ve} = +ve\]

OpenStudy (vishweshshrimali5):

got this ?

OpenStudy (anonymous):

Yes

OpenStudy (vishweshshrimali5):

ok gud \[\frac{\frac{-2}{3}}{-18}=\frac{\frac{2}{3}}{18}\]

OpenStudy (anonymous):

alright whats next

OpenStudy (vishweshshrimali5):

ok very gud now \[\frac{\frac{2}{3}}{18} = \frac{2}{3*18}\]

OpenStudy (anonymous):

its 2/54

OpenStudy (vishweshshrimali5):

yes \[WONDERFUL\]

OpenStudy (anonymous):

THats it?

OpenStudy (vishweshshrimali5):

yes \[\frac{\frac{-2}{3}}{18} = \frac{2}{54} = \frac{1}{27}\]

OpenStudy (anonymous):

Thank You, you earned yourself a medal

OpenStudy (vishweshshrimali5):

\[THANKS\]